Qualitative Analysis Of Two Dimensional System Differential Equations Of Special Form

Yahyo Muhtorov Samarkand state University

The whole system of two differential equations, right parts which consists of equation of m grade that corresponds to the equation of the searching function with gradual coefficients into which changeable enter to m grade (m.n. the natural numbers) We got enough conditions about absence of limit cycles. In dependence on legibility searching allocation of the special points of the system.